OpenStudy (anonymous):

15=-2(2t-1) help plzzz

OpenStudy (anonymous):

15 = 2(2t-1) 4t - 2 = 15 4t = 17 t = 17/4

OpenStudy (anonymous):

15=-4t+2 t=17/4